OpenStudy (anonymous):

Factor this expression. a2-6a-7 ... need help >.<

OpenStudy (akshay_budhkar):

\[a^2-7a+a-7=a(a-7)+1(a-7)=(a-7)(a+1)\]

OpenStudy (anonymous):

thankies ^_^

OpenStudy (akshay_budhkar):

yw :D